OpenStudy (anonymous):

Please explain. What would you do if you didn’t know what the values of the denominators were, such as when there are variables in the denominators? For example, what denominator would you use if you were adding 1/a and 1/b ? How would your new denominator affect the numerators?

OpenStudy (emmamink):

You would not use one because variables stand for unknown values. If you have no way to figure that out, then you would just keep 1/a + 1/b as it is.

OpenStudy (emmamink):

Right @TheSmartOne

OpenStudy (emmamink):

Now, if that answered that question, feel free to medal, and if it didn't message me and let me know so I can try harder in the future to help you.

OpenStudy (anonymous):

Thank you @EmmaMink

OpenStudy (emmamink):

Your welcome!

TheSmartOne (thesmartone):

\(\Huge \frac{1}{a}+\frac{1}{b}=\frac{a+b}{ab}\)

TheSmartOne (thesmartone):

So you can change it..

OpenStudy (anonymous):

please explain @TheSmartOne

TheSmartOne (thesmartone):